OpenStudy (anonymous):

Write the standard form of the line that contains a slope of 2/3 and passes through the point (1, 1).

Nnesha (nnesha):

\[\huge\rm y = mx + b \] slope intercept form where m is slope and b is y-intercept so replace m by 2/3 and x and y by (1,1) then solve for b

OpenStudy (anonymous):

How?

Nnesha (nnesha):

first replace m and x y values and let me know what you get after that

Nnesha (nnesha):

by**

OpenStudy (anonymous):

actually you can use the slope intercept form: y - y0 = m(x-x0)

OpenStudy (anonymous):

where the coordinates are (x0, y0) and the slope is m

Nnesha (nnesha):

^^that one is point slope form :)

OpenStudy (anonymous):

@Kommander_Kitten

OpenStudy (anonymous):

@AlexandervonHumboldt2

Nnesha (nnesha):

replace m by 2/3 x by 1 y by 1 in order pair first number is x and 2nd one is y \[\huge\rm (x ,y )\] do it and show ur work

OpenStudy (anonymous):

I don't understand

Nnesha (nnesha):

what you don't understand ?

OpenStudy (anonymous):

What I'm supposed to do

Nnesha (nnesha):

you suppose to replace m by 2/3 x by 1 and y by 1 into this equation y=mx + b

OpenStudy (anonymous):

But it's supposed to be in standard form, not slope-intercept

Nnesha (nnesha):

ik first you have to solve for b(y-intercept)

OpenStudy (anonymous):

Okay so 1=2/3(1)+b

Nnesha (nnesha):

yes solve for b

OpenStudy (anonymous):

How..lol

OpenStudy (anonymous):

b=1/3

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

So now what?

Nnesha (nnesha):

standard form is Ax +By = C

OpenStudy (anonymous):

OMG I feel like such a blonde now -_- lol

OpenStudy (anonymous):

2/3x+1/3y=1 ..?

Nnesha (nnesha):

the equation u just got is y = 2/3x + 1/3 right both have same denominator so we can take it common \[\frac{ 2 }{ 3 }x +\frac{ 1 }{ 3 } = \frac{ 2x +1 }{ 3 }\] so we can wirte like this

Nnesha (nnesha):

write*

Nnesha (nnesha):

\[\huge\rm y = \frac{ 2x +1 }{ 3 }\] multiply both side by 3

OpenStudy (anonymous):

6x+6/9 ?

Nnesha (nnesha):

nope

Nnesha (nnesha):

 both side multiply by 3